OpenStudy (dinorap1):

WILL GIVE MEDAL FOR BEST ANSWER! Solve the system by elimination. {-2x+2y+3z=0 {-2x-y+z=-3 {2x+3y+3z=5

OpenStudy (compassionate):

Super simple stuff. Just solve two, then solve the last one by plugging in your x-y values

OpenStudy (dinorap1):

How do I start? I have a general idea of how to do this, I'm just confused about a couple of things.

OpenStudy (dinorap1):

By the way, imagine that they are all in one big bracket. That's how it's supposed to be, but I don't know how to do that on here.

OpenStudy (jdoe0001):

hmmm have you done system of equations yet? with 2 variables? how abotu with 3 variables?

OpenStudy (dinorap1):

I have done a system of equations before, with 3 variables and 2 variables. I'm a little rusty though, and don't really remember how to start...

OpenStudy (compassionate):

-2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 We need to eliminate our first equation. We can do this by eliminating z. Multiply out second equation by -3. -2x + 2y + 3z = 0 6x + 3y - 3z = 9 2x + 3y + 3z = 5 We we can eliminate z by adding 4x + 5y = 9 2x + 3y + 3z = 5 So now we do the same thing for 2x + 3y + 3z = 5. Eliminate it and remove z, then you'll hve a system with two (,x, y) points you can solve.

OpenStudy (jdoe0001):

well... pick any two equations from the three available ones and cancel out 1 of the variables

OpenStudy (jdoe0001):

then you pick another set of 2 off the three available and cancel out the SAME variable then use the two RESULTANT equations from each cancellation and solve it, like you would any system of equations of two variables

OpenStudy (dinorap1):

Sooo I think the way you're supposed to do it is to add/subtract 2 equations from each other, and then use the result of those 2 equations to solve for one or 2 variables, right? Is that the way you guys are doing it?

OpenStudy (compassionate):

After that, plug in both your x and y values to find z. Tada!

OpenStudy (dinorap1):

So you cancel out z first, right? I'm a little confused about the way you wrote it...

OpenStudy (dinorap1):

You multiplied the second equation by -3 and then subtracted that equation from both of the other equations to cancel out the z?

OpenStudy (mathmate):

(1)+(3).......(4) (2)+(3).......(5) (4) and (5) contain only y and z, i.e. 2x2.

OpenStudy (dinorap1):

@mathmate Ohh the way you wrote it is much easier to understand, no offense. Okay, I'll try it your way.

OpenStudy (mathmate):

It is quite true that eliminating any variable will work. Sometimes you will notice that certain variables are easier to be eliminated than others! :)

OpenStudy (dinorap1):

@mathlete Alright, so equations (4) and (5) are: {5y+6z=5 {2y+4z=2 Sooo... now what? Now do you just solve them like a two variable equation? I kinda forgot how to do this a little...

OpenStudy (dinorap1):

@mathmate I got z=0. Is that wrong? It seems wrong...

OpenStudy (mathmate):

z=0 is correct, but you still have to find the others.

OpenStudy (dinorap1):

Okay, good! I was just wondering if I did it right. So how do I find the others now?

OpenStudy (mathmate):

Choose any two of the equation (say the first two), put z=0 and solve again a system of two unknowns (x and y).

OpenStudy (dinorap1):

@mathmate What 2 equations could I use? The original equations, or equations (4) and (5)?

OpenStudy (mathmate):

Equation (4) and (5) do not contain x any more, so you must use two of (1), (2) or (3).

OpenStudy (dinorap1):

Okay, but could I plug (z=0) into (4) or (5) to get the value for y?

OpenStudy (dinorap1):

@mathmate

OpenStudy (mathmate):

Yes, that's even better, lol!

OpenStudy (dinorap1):

Yay! I think I can figure out the rest then. Thank you so much for all your help! :)

OpenStudy (mathmate):

Good! You're welcome! :)

OpenStudy (mathmate):

What did you get for the final answer? (?,?,?)

OpenStudy (dinorap1):

@mathmate I got (1,1,0) I'm about to test it out to see if I'm right...

OpenStudy (dinorap1):

Yep, it's the correct answer! Yaaay! Thank you so much for helping! This was giving me such a headache... phew! Glad it's finally solved! :)

OpenStudy (anonymous):

hey could someone help me with my math cause i'm really confused and i need someone to explain

OpenStudy (dinorap1):

@athalia I'm sure somebody will help you... good luck with it! It looks too confusing for me! Sorry!

OpenStudy (anonymous):

wow ok thxs anyway

OpenStudy (mathmate):

@dinorap1 Good, that was the correct answer, guess you have already checked!